# SpiderCat: A Catalog of Compact Binary Millisecond Pulsars

### v1

@@ 3 Results and discussion
In this section, we present key statistical and physical properties of the 109 compact binary millisecond pulsars compiled in SpiderCat, focusing on their Galactic distribution, spin behavior, orbital characteristics, companion masses, and distances. Through multi-wavelength data and population-wide trends, we provide insight into the demographics, evolution, and spatial distribution of spider MSPs. The following subsections highlight our main findings derived from the catalog, but independent analysis using the public SpiderCat database may uncover more emerging properties.
@@ 3.1 Galactic distribution
Figure 2: Galactic distribution of spider systems in a Mollweide projection, color-coded by spider subclass. The top and left panels show histograms of Galactic longitude (bin size: 12.4) and latitude (bin size: 9.5), respectively. The spider population is strongly concentrated toward low Galactic latitudes, while the distribution in Galactic longitude is relatively uniform.
Figure 3: Histogram of estimated Galactic heights for spider systems (bin size: 0.33 kpc), color-coded by spider subclass. The distribution is approximately normal, with a mean height of $\bar{z}=0.13\pm 0.04$ kpc and a standard deviation of $\sigma_{z}=0.97\pm 0.07$ kpc, overplotted as a scaled Gaussian (dashed line).
The Galactic distribution of spiders is shown in Fig. 2. Their Galactic latitude distribution is concentrated toward low latitudes, with the majority of sources located at $|b|<30^{\circ}$ (73%), including 27 sources (25%) within $|b|<6^{\circ}$. The distribution in Galactic longitude is relatively uniform, with most sources (68%) found at $|l|<70^{\circ}$. There appears to be a concentration between $20^{\circ}<l<70^{\circ}$, where 35 sources (32%) are located. However, we note that this region includes 12 sources from FAST-GPPS, which may introduce an observational bias. Most spider searches to date have avoided the Galactic plane ($|b|>5^{\circ}$), to minimize $\gamma$-ray diffuse emission and/or reddening. On the other hand, the intrinsic number density of spiders (like that of most stellar populations) is expected to increase when approaching the Galactic plane.
@@ 3.2 Spin periods
Figure 4: Spin period ($P$) versus spin period derivative ($\dot{P}$) diagram for ATNF pulsars, highlighting binaries and confirmed spider systems. Lines of constant $\dot{E}$, $B$ and $\tau$ are superposed. Also, the classical and outer magnetosphere “death lines” are shown as dashed blue and dash-dotted green lines, respectively (see text). Histograms of $P$ and $\dot{P}$ are shown above and to the right of the plot, respectively, on a logarithmic scale for clarity. Spiders occupy the lower-left region of the $P/\dot{P}$-diagram, corresponding to the fastest-spinning, most weakly magnetized, and oldest pulsars. Two outlier systems, PSRs J2129--0429 and J1932$+$2121, lie above the main population, likely representing mildly recycled systems with longer spin periods and higher $\dot{P}$, indicative of alternative evolutionary paths.
@@
Pulsars found in spider systems are among the fastest-spinning MSPs, including the record holder BW PSR J0952--0607 (L. Nieder et al., 2019), which has a spin period of just 1.41 ms (the fastest after the globular cluster RB PSR J1748–2446ad, J. W. T. Hessels et al. 2006). In the $P-\dot{P}$-diagram (Figure 4), spiders occupy the lower-left corner at the end of the recycled pulsar branch and are characterized by short spin periods (median of 2.5 ms, with nearly all systems falling in the narrow 1.5–5.0 ms range) and spin period derivatives around $\dot{P}\approx 10^{-20}$ s/s. As a result, their characteristic ages ($\tau\sim P/2\dot{P}$) are between $10^{9}$ and $10^{10}$ years. Their characteristic surface magnetic fields ($B\propto\sqrt{P\dot{P}}$) are among the lowest inferred for pulsars, between $10^{8}$ and $10^{9}$ G. The spin-down luminosities ($\dot{E}\propto\dot{P}/P^{3}$) of spider systems range from $4\times 10^{33}$ erg/s to $3\times 10^{35}$ erg/s (see Table 2.2).
@@
We also show two theoretical “death lines” in Fig. 4, which represent approximate thresholds below which pair production, and hence pulsar activity, is expected to cease (P. A. Sturrock, 1971; M. A. Ruderman & P. G. Sutherland, 1975; K. Chen & M. Ruderman, 1993). The classical death line, derived under the polar cap particle acceleration model in a dipole magnetic field, is shown as a blue dashed line and corresponds to [^6] Here we have converted the original relation between surface magnetic field strength and spin period into a relation between spin period derivative and spin period by assuming $B[\rm{G}]=\sqrt{\frac{3c^{3}I\dot{P}P}{8\pi^{2}R^{6}}}\approx 3.2\times 10^{19}\sqrt{P\dot{P}}$, where $I=10^{45}$ g cm2 and $R=10$ km. $\log{\dot{P}}=2.75\times P-14.35$. An alternative outer magnetosphere death line, relevant for pulsars with high spin-down power, is shown as a green dash-dotted line and corresponds to $\log{\dot{P}}=3.8\times P-11.16$. All spider systems lie above both death lines, with the exception of PSR J1932$+$2121, suggesting it may be near the limit for sustained $\gamma$-ray production.
@@
Two notable outliers in the $P-\dot{P}$-diagram are PSR J2129--0429 ($P=7.5$ ms, $\dot{P}=3\times 10^{-19}$ s/s) and PSR J1932$+$2121 ($P=14.2$ ms, $\dot{P}=3.5\times 10^{-19}$ s/s), both of which have longer $P$ and higher $\dot{P}$ than the bulk of spiders. These sources likely represent mildly recycled systems, possibly due to an inefficient mass-transfer phase or an early stage of evolutionary development, after which a subsequent Roche-Lobe overflow phase may further spin-up the pulsar (D. Misra et al., 2025b).
@@ 3.4 Distances
Measuring distances to spider systems is challenging, as they are typically inferred from the dispersion measure (DM) of the radio pulsar and require models of the free electron density along the line of sight. In this paper, we estimate distances to spiders either via the DM-based method with the Galactic electron density model of J. M. Yao et al. (2017, YMW16) where available [^7] Using the pygedm package (D. C. Price et al., 2021)., or, alternatively, through geometric distances derived from Gaia DR3 parallaxes. For the latter, we adopt the distance prior from C. A. L. Bailer-Jones et al. (2021), and use the parallax-distance estimate even in cases where the parallax error-to-signal ratio exceeds 1 (when the distance posterior is heavily influenced by the prior). See also K. I. I. Koljonen & M. Linares (2023, Appendix A) for details on distance estimates based on Gaia DR3 parallaxes.
@@
$\bullet$ PSR J1720-0533: We adopt the DM distance using the Galactic electron density model of J. M. Cordes & T. J. W. Lazio (2002, NE2001). As discussed in K. I. I. Koljonen et al. (2024), the inclusion of a dense component toward the North Polar Spur in the YMW16 model underestimates the distance to this BW, as revealed by flux upper limits on its IR counterpart.
$\bullet$ PSR J0955-3947 and PSR J1036-4353: We use NE2001 DM distances, which are more consistent with Gaia parallax estimates (see comparison in K. I. I. Koljonen & M. Linares 2023).
$\bullet$ PSR J1317-0157: We adopt the NE2001 DM distance, as the YMW16 model yields a pegged and likely unphysical distance of 25 kpc.
$\bullet$ PSR J0838-2827: We use the distance derived from Gaia parallax measurements (see discussion in T. Thongmeearkom et al. 2024).
$\bullet$ PSR J0636+5128: We rely on the NE2001 DM distance (see discussion in P. Draghis & R. W. Romani 2018).
@@ 3.5 Multiwavelength counterparts
We first searched for matches within 1 arcmin of our spider locations in the ATNF Pulsar Catalog (R. N. Manchester et al., 2005, v. 2.6.0), using psrqpy (M. Pitkin, 2018, v. 1.3.2). We used the following observational catalogs to search for associated multiwavelength counterparts to the spiders listed in Table 2.2: Gaia Data Release 3 (Gaia DR3; Gaia Collaboration et al., 2016, 2022), the Two Micron All Sky Survey (2MASS; M. F. Skrutskie et al., 2006), Panoramic Survey Telescope and Rapid Response System Data Release 2 (PS1 DR2; K. C. Chambers et al., 2016), SkyMapper Southern Sky Survey Data Release 4 (SMSS DR4; C. A. Onken et al., 2024), the Swift/XRT Point Source Catalogue (2SXPS; P. A. Evans et al., 2020), Chandra Source Catalog Release 2.1 (CSC2.1; I. N. Evans et al., 2024), eROSITA-DE Data Release 1 (eRASS1; A. Merloni et al., 2024), the XMM-Newton Serendipitous Source Catalogue Data Release 14 (4XMM-DR14; N. A. Webb et al., 2020) and the Fermi-LAT 14-year Source Catalog (4FGL-DR4; S. Abdollahi et al., 2022b; J. Ballet et al., 2023). We applied a search radius of 2 arcsec for optical/IR catalogs, 4 arcsec for X-ray catalogs, and 10 arcmin for the $\gamma$-ray catalog, centered on the positions (RA and Dec) listed in Table 2.2. Source IDs of the identified counterparts are presented in Tables 3 and 4, along with their angular separations from the query coordinates. Table 5 includes distance estimates, X-ray and $\gamma$-ray fluxes and luminosities, as well as g-band apparent and absolute magnitudes, and g–r color indices of the counterparts.
@@ 3.5.1 X-rays and $\gamma$-rays
Table 3: X-ray/$\gamma$-ray counterparts. See Section 3.5.1 for details.
Name Fermi sep XMM-Newton sep Swift sep Chandra sep eROSITA sep
4FGL (’) 4XMM (”) 2SXPS (”) 2CXO (”) 1eRASS (”)
PSR J0023+0923 J0023.4+0920 3.1 J002316.8+092323 0.1
PSR J0212+5321 J0212.1+5321 0.6 159668 0.6 J021210.4+532138 0.0
@@
PSR J2241-5236 J2241.7-5236 0.5 J224142.0-523635 0.4 167739 3.0 J224142.0-523636 0.5
PSR J2256-1024 J2256.8-1024 1.1 J225656.3-102434 0.3
PSR J2333-5526 J2333.1-5527 1.8 J233316.0-552620 0.5
PSR J2339-0533 J2339.6-0533 0.1 J233938.7-053305 0.3 54830 1.1 J233938.7-053305 0.2
Figure 6: The relationship between X-ray ($L_{\rm X}$) and $\gamma$-ray ($L_{\gamma}$) luminosities for spider systems. $\gamma$-ray luminosities are not significantly different between spider types. On the other hand, RBs exhibit systematically higher X-ray luminosities likely due to stronger intrabinary shocks. BWs follow a relation of $L_{\rm X}\propto L_{\gamma}^{0.6\pm 0.2}$, while RBs may follow a steeper trend of $L_{\rm X}\propto L_{\gamma}^{1.0\pm 0.3}$, suggesting differing emission contributions across populations.
Figure 7: Histogram of the logarithm of the ratio of $\gamma$-ray ($F_{\gamma}$) to X-ray ($F_{\rm X}$) fluxes for spider systems (bin size: 0.25 dex), color-coded by spider type. The flux ratios for BWs are more narrowly distributed and generally higher. For $\log\,(F_{\gamma}/F_{\rm X})<2.3$, it becomes increasingly likely that the system is a RB.
We associate a total of 84 Fermi 4FGL sources and 55 X-ray counterparts with spider systems (see Table 1), corresponding to 77% and 50% of the catalog, respectively. Some sources appear in earlier Fermi catalogs (2FGL J0846.0$+$2820; S. J. Swihart et al. 2017, and possibly PSR J1723--2837; C. Y. Hui et al. 2014), or in Fermi 3PC catalog (PSR J0630$+$5128; D. A. Smith et al. 2023) but are absent from the 4FGL catalog, suggesting intermittent and/or weak $\gamma$-ray emission. Including these three sources raises the $\gamma$-ray association rate among spiders to 80%.
Twenty-one sources (19%) lack high-energy counterparts. Of these, 12 originate from the FAST-GPPS survey (P. F. Wang et al., 2025) and suffer from low-precision radio positions and high Galactic background. Three are BWs with very large DM-inferred distances (PSR J2055$+$3829, PSR J1928$+$1245, PSR J1745--23) and likely lie beyond the sensitivity limits of current surveys. PSR J1720--0533, also a BW without high-energy detection, is probably located farther away ($\sim$3 kpc) than suggested by its DM-derived distance (K. I. I. Koljonen et al., 2024). Two sources are spider candidates based on optical variability (3FGL J2221.6$+$6507 and ZTF J1406$+$1222). This leaves three sources with moderate distances ($\lesssim$2.5 kpc), two BWs (PSR J1602--1009 and PSR J1630$+$3550) and one RB (PSR J1242--4712), for which deeper high-energy counterpart searches are encouraged.
We computed X-ray and $\gamma$-ray luminosities using the X-ray fluxes from the aforementioned catalogs and the distances described in Section 3.4. The $\gamma$-ray energy fluxes are quoted in the Fermi-LAT 0.1 – 100 GeV energy band, obtained by spectral fitting (as tabulated in 4FGL-DR4). For X-ray luminosity estimates, fluxes were used in the following priority order: XMM-Newton (0.2–12 keV), Neil Gehrels Swift Observatory (0.3–10 keV), Chandra X-ray Observatory (0.2–7 keV), and eROSITA onboard Spektrum-Roentgen-Gamma (0.2–8 keV). The resulting luminosities are provided in Table 5.
Figure 6 displays the X-ray and $\gamma$-ray luminosities of the spider systems. The $\gamma$-ray luminosities span $10^{32}$$10^{35}$ erg/s with a mean luminosity between $10^{33}$ and $10^{33.5}$ erg/s. No significant differences are observed between spider subtypes in their $\gamma$-ray luminosities.
Spider X-ray luminosities, known to exhibit a bimodal distribution (J. Lee et al., 2018; S. J. Swihart et al., 2022; K. I. I. Koljonen & M. Linares, 2023), range from $10^{30}$ to $10^{34}$ erg/s. RBs consistently display higher X-ray luminosities, with values exceeding $10^{32}$ erg/s found exclusively among them (see Fig. 6). This enhanced emission likely originates from stronger intrabinary shocks (e.g., R. W. Romani & N. Sanchez, 2016; Z. Wadiasingh et al., 2017, 2018; C. J. T. van der Merwe et al., 2020).
@@
In Figure 7, we present the histogram of the logarithm of the $\gamma$-ray to X-ray flux (or luminosity) ratios. BWs exhibit a much narrower distribution, with a mean value of $\log\,(F_{\gamma}/F_{\rm X})=2.6\pm 0.25$, whereas RBs show a broader distribution, centered at $\log\,(F_{\gamma}/F_{\rm X})=1.6\pm 0.8$. The lowest observed ratio for BWs is $\log\,(F_{\gamma}/F_{\rm X})\approx 2.3$; sources with lower values are thus likely to be RBs. Therefore, the flux ratio can serve as a diagnostic tool for distinguishing between source types when both X-ray and $\gamma$-ray fluxes are available. This represents an update to the flux ratios reported by D. Salvetti et al. (2017), in which such distinctions could not be clearly identified.
3.5.2 Optical/IR
@@
Table 4: Optical/infrared counterparts. See Section 3.5.2 for details.
Name Gaia sep Pan-STARRS sep SkyMapper sep 2MASS sep

### v4

@@ 3 Results and Discussion
In this section, we present key statistical and physical properties of the 109 compact binary millisecond pulsars compiled in SpiderCat, focusing on their Galactic distribution, spin behavior, orbital characteristics, companion masses, and distances. Through multiwavelength data and population-wide trends, we provide insight into the demographics, evolution, and spatial distribution of spider MSPs. The following subsections highlight our main findings derived from the catalog, but independent analysis using the public SpiderCat database may uncover more emerging properties.
@@ 3.1 Galactic Distribution
Figure 2: Galactic distribution of spider systems in a Mollweide projection, color coded by spider subclass. The top and left panels show histograms of Galactic longitude (bin size: 12.4) and latitude (bin size: 9.5), respectively. The spider population is strongly concentrated toward low Galactic latitudes, while the distribution in Galactic longitude is relatively uniform.
Figure 3: Histogram of estimated Galactic heights for spider systems (bin size: 0.33 kpc), color coded by spider subclass. The distribution is approximately normal, with a mean height of $\bar{z}=0.14\pm 0.09$ kpc and a standard deviation of $\sigma_{z}=0.93\pm 0.06$ kpc, overplotted as a scaled Gaussian (dashed line).
The Galactic distribution of spiders is shown in Figure 2. Their Galactic latitude distribution is concentrated toward low latitudes, with the majority of sources located at $|b|<30^{\circ}$ (73%), including 27 sources (25%) within $|b|<6^{\circ}$. The distribution in Galactic longitude is relatively uniform, with most sources (68%) found at $|l|<70^{\circ}$. There appears to be a concentration between $20^{\circ}<l<70^{\circ}$, where 35 sources (32%) are located. However, we note that this region includes 12 sources from FAST-GPPS, which may introduce an observational bias. Most spider searches to date have avoided the Galactic plane ($|b|>5^{\circ}$), to minimize $\gamma$-ray diffuse emission and/or reddening. On the other hand, the intrinsic number density of spiders (like that of most stellar populations) is expected to increase when approaching the Galactic plane.
@@ 3.2 Spin Periods
Figure 4: Spin period ($P$) vs. spin period derivative ($\dot{P}$) diagram for ATNF pulsars, highlighting binaries and confirmed spider systems. Lines of constant $\dot{E}$, $B$ and $\tau$ are superposed. Also, the classical and outer magnetosphere “death lines” are shown as dashed blue and dashed-dotted green lines, respectively (see text). Histograms of $P$ and $\dot{P}$ are shown above and to the right of the plot, respectively, on a logarithmic scale for clarity. Spiders occupy the lower-left region of the $P/\dot{P}$-diagram, corresponding to the fastest-spinning, most weakly magnetized, and oldest pulsars. Two outlier systems, PSRs J2129-0429 and J1932++2121, lie above the main population, likely representing mildly recycled systems with longer spin periods and higher $\dot{P}$, indicative of alternative evolutionary paths.
@@
Pulsars found in spider systems are among the fastest-spinning MSPs, including the record holder BW PSR J0952-0607 (L. Nieder et al., 2019), which has a spin period of just 1.41 ms (the fastest after the globular cluster RB PSR J1748–2446ad, J. W. T. Hessels et al. 2006). In the $P-\dot{P}$-diagram (Figure 4), spiders occupy the lower-left corner at the end of the recycled pulsar branch and are characterized by short spin periods (median of 2.5 ms, with nearly all systems falling in the narrow 1.5–5.0 ms range) and spin period derivatives around $\dot{P}\approx 10^{-20}$ s s-1. As a result, their characteristic ages ($\tau\sim P/2\dot{P}$) are between $10^{9}$ and $10^{10}$ yr. Their characteristic surface magnetic fields ($B\propto\sqrt{P\dot{P}}$) are among the lowest inferred for pulsars, between $10^{8}$ and $10^{9}$ G. The spin-down luminosities ($\dot{E}\propto\dot{P}/P^{3}$) of spider systems range from $4\times 10^{33}$ erg s-1 to $3\times 10^{35}$ erg s-1 (see Table 2).
@@
We also show two theoretical “death lines” in Figure 4, which represent approximate thresholds below which pair production, and hence pulsar activity, is expected to cease (P. A. Sturrock, 1971; M. A. Ruderman & P. G. Sutherland, 1975; K. Chen & M. Ruderman, 1993). The classical death line, derived under the polar cap particle acceleration model in a dipole magnetic field, is shown as a blue dashed line and corresponds to [^7] Here we have converted the original relation between surface magnetic field strength and spin period into a relation between spin period derivative and spin period by assuming $B[\rm{G}]=\sqrt{\frac{3c^{3}I\dot{P}P}{8\pi^{2}R^{6}}}\approx 3.2\times 10^{19}\sqrt{P\dot{P}}$, where $I=10^{45}$ g cm2 and $R=10$ km. $\log{\dot{P}}=2.75\times P-14.35$ (corresponding to Equation (6) in K. Chen & M. Ruderman 1993).
For an outer magnetosphere accelerator model, which is relevant for pulsars with high spin-down power (K. S. Cheng et al., 1986a, b), an alternative death line marking the cessation of pair production and strong $\gamma$-ray emission in the outer magnetosphere is shown as a green dash-dotted line and corresponds to $\log{\dot{P}}=3.8\times P-11.16$ (Equation (27) in K. Chen & M. Ruderman 1993). All spider systems lie above both death lines, with the exception of PSR J1932++2121, suggesting it may be near the limit for sustained $\gamma$-ray production.
@@
Two notable outliers in the $P-\dot{P}$-diagram are PSR J2129-0429 ($P=7.5$ ms, $\dot{P}=3\times 10^{-19}$ s s-1) and PSR J1932++2121 ($P=14.2$ ms, $\dot{P}=3.5\times 10^{-19}$ s s-1), both of which have longer $P$ and higher $\dot{P}$ than the bulk of spiders. Notably, both are classified as RBs, which may have spent less time in the spider phase compared to typical BWs. This may account for their relatively mild recycling, potentially resulting from an inefficient mass-transfer phase or an early evolutionary stage, with the possibility of further spin-up during a subsequent Roche-Lobe overflow phase (D. Misra et al., 2025a).
A limited amount of accretion would, in turn, imply low pulsar masses. Indeed, evolutionary calculations presented in D. Misra et al. (2025a) support neutron star masses of $M_{\rm NS}\approx 1.6\,M_{\odot}$. While no mass estimate currently exists for PSR J1932++2121, C. J. Clark et al. (2023a) has constrained the mass of PSR J2129-0429 to $1.61\,M_{\odot}<M_{\rm NS}<1.88\,M_{\odot}$ (95% confidence).
@@ 3.4 Distances
Measuring distances to spider systems is challenging, as they are typically inferred from the DM of the radio pulsar and require models of the free electron density along the line of sight. In this paper, we estimate distances to spiders either via the DM-based method with the Galactic electron density model of J. M. Yao et al. (2017, “YMW16”) where available, [^8] Using the pygedm package (D. C. Price et al., 2021). or, alternatively, through geometric distances derived from Gaia DR3 parallaxes. In cases where the relative parallax uncertainty ($\sigma_{\varpi}/\varpi$) is less than 0.2, we always adopt the parallax-based distance, regardless of whether a DM-based estimate is available. In deriving the parallax-based distance we adopt the prior from C. A. L. Bailer-Jones et al. (2021), and use the parallax-distance estimate even in cases where the parallax error-to-signal ratio exceeds one (when the distance posterior is heavily influenced by the prior). See also K. I. I. Koljonen & M. Linares (2023, Appendix A) for details on distance estimates based on Gaia DR3 parallaxes.
@@
  1. 1.
    PSR J1720-0533. We adopt the DM distance using the Galactic electron density model of J. M. Cordes & T. J. W. Lazio (2002, “NE2001”). As discussed in K. I. I. Koljonen et al. (2024), the inclusion of a dense component toward the North Polar Spur in the YMW16 model underestimates the distance to this BW, as revealed by flux upper limits on its IR counterpart.
  2. 2.
    PSR J0955-3947 and PSR J1036-4353. We use NE2001 DM distances, which are more consistent with Gaia parallax estimates (see comparison in K. I. I. Koljonen & M. Linares 2023).
  3. 3.
    PSR J1317-0157. We adopt the NE2001 DM distance, as the YMW16 model yields a pegged and likely unphysical distance of 25 kpc.
  4. 4.
    PSR J0838-2827. We use the distance derived from Gaia parallax measurements (see discussion in T. Thongmeearkom et al. 2024).
  5. 5.
    PSR J0636+5128. We rely on the NE2001 DM distance (see discussion in P. Draghis & R. W. Romani 2018).
@@ 3.5 Multiwavelength Counterparts
[…]
@@ 3.5.1 X-Rays and $\gamma$-Rays
Table 3: X-Ray and $\gamma$-Ray Counterparts footnotetext: Notes. See Section 3.5.1 for details. (This table is available in machine-readable form in the online article.)
Name Fermi Sep. XMM-Newton Sep. Swift Sep. Chandra Sep. eROSITA Sep.
[…]
PSR J0023+0923 J0023.4+0920 3.1 J002316.8+092323 0.12
PSR J0212+5321 J0212.1+5321 0.6 159668 0.6 J021210.4+532138 0.04
@@
PSR J2241-5236 J2241.7-5236 0.5 J224142.0-523635 0.42 167739 3.0 J224142.0-523636 0.46
PSR J2256-1024 J2256.8-1024 1.1 J225656.3-102434 0.34
PSR J2333-5526 J2333.1-5527 1.8 J233316.0-552620 0.45 143919 6.0
PSR J2339-0533 J2339.6-0533 0.1 J233938.7-053305 0.27 54830 1.1 J233938.7-053305 0.2
Figure 6: The relationship between X-ray ($L_{\rm X}$) and $\gamma$-ray ($L_{\gamma}$) luminosities for spider systems. $\gamma$-ray luminosities are not significantly different between spider types. On the other hand, RBs exhibit systematically higher X-ray luminosities likely due to stronger IBSs. BWs follow a relation of $L_{\rm X}\propto L_{\gamma}^{0.5\pm 0.2}$, while RBs may follow a steeper trend of $L_{\rm X}\propto L_{\gamma}^{0.9\pm 0.3}$, suggesting differing emission contributions across populations.
Figure 7: Histogram of the logarithm of the ratio of $\gamma$-ray ($F_{\gamma}$) to X-ray ($F_{\rm X}$) fluxes for spider systems (bin size: 0.25 dex), color coded by spider type. The flux ratios for BWs are more narrowly distributed and generally higher. For $\log\,(F_{\gamma}/F_{\rm X})<2$, it becomes increasingly likely that the system is an RB.
@@
We associate a total of 86 Fermi 4FGL sources and 55 X-ray counterparts with spider systems (see Table 1), corresponding to 77% and 50% of the catalog, respectively. Some sources appear in earlier Fermi catalogs (2FGL J0846.0++2820, S. J. Swihart et al. 2017; and possibly PSR J1723-2837, C. Y. Hui et al. 2014), or in Fermi 3PC catalog (PSR J0630++5128, D. A. Smith et al. 2023) but are absent from the 4FGL catalog, suggesting intermittent and/or weak $\gamma$-ray emission. Including these three sources raises the $\gamma$-ray association rate among spiders to 80%.
Twenty-two sources (20%) lack high-energy counterparts. Of these, 12 originate from the FAST-GPPS survey (P. F. Wang et al., 2025) and suffer from low-precision radio positions and high Galactic background. Three are BWs with very large DM-inferred distances (PSR J2055++3829, PSR J1928++1245, and PSR J1745-23), and one is an RB, PSR J1932++2121, which has a relatively large Gaia parallax-based distance estimate ($\sim$5 kpc). These characteristics likely place them beyond the sensitivity limits of current surveys. PSR J1720-0533, also a BW without high-energy detection, is probably located farther away ($\sim$3 kpc) than suggested by its DM-derived distance (K. I. I. Koljonen et al., 2024). Two sources are spider candidates based on optical variability (3FGL J2221.6++6507 and ZTF J1406++1222). This leaves three sources with moderate distances ($\lesssim$2.5 kpc), two BWs (PSR J1602-1009 and PSR J1630++3550) and one RB (PSR J1242-4712), which represent promising targets for deeper X-ray and $\gamma$-ray follow-up observations aimed at identifying potential high-energy counterparts.
We computed X-ray and $\gamma$-ray luminosities using the X-ray fluxes from the aforementioned catalogs and the distances described in Section 3.4. The $\gamma$-ray energy fluxes are quoted in the Fermi-LAT 0.1–100 GeV energy band, obtained by spectral fitting (as tabulated in 4FGL-DR4). For X-ray luminosity estimates, fluxes were used in the following priority order: XMM-Newton (0.2–12 keV), Neil Gehrels Swift Observatory (0.3–10 keV), Chandra X-ray Observatory (0.2–7keV), and eROSITA onboard Spektrum-Roentgen-Gamma (0.2–8 keV). The resulting luminosities and the observatory used in each case are provided in Table 5.
Figure 6 displays the X-ray and $\gamma$-ray luminosities of the spider systems. The $\gamma$-ray luminosities span $10^{32}$$10^{35}$ erg s-1 with a mean luminosity between $10^{33}$ and $10^{33.5}$ erg s-1. No significant differences are observed between spider subtypes in their $\gamma$-ray luminosities.
Spider X-ray luminosities, known to exhibit a bimodal distribution (J. Lee et al., 2018; S. J. Swihart et al., 2022; K. I. I. Koljonen & M. Linares, 2023), range from $10^{30}$ to $10^{34}$ erg s-1. RBs consistently display higher X-ray luminosities, with values exceeding $10^{32}$ erg s-1 found exclusively among them (see Figure 6). This enhanced emission likely originates from stronger IBSs (e.g., R. W. Romani & N. Sanchez, 2016; Z. Wadiasingh et al., 2017, 2018; C. J. T. van der Merwe et al., 2020). Additionally, in RBs, the IBS tends to wrap around the pulsar rather than the companion (Z. Wadiasingh et al., 2017; D. Kandel et al., 2019; A. G. Sullivan & R. W. Romani, 2025). This geometry places the shock closer to the pulsar, resulting in a higher incident particle flux and more efficient conversion of spin-down power into X-rays, further contributing to the observed luminosity differences between the two populations.
@@
In Figure 7, we present a histogram of the logarithm of the $\gamma$-ray to X-ray flux (or luminosity) ratios. BWs exhibit a much narrower distribution, with a mean value of $\log\,(F_{\gamma}/F_{\rm X})=2.6\pm 0.3$, whereas RBs show a broader distribution, centered at $\log\,(F_{\gamma}/F_{\rm X})=1.6\pm 0.8$. The lowest observed ratio for BWs is $\log\,(F_{\gamma}/F_{\rm X})\approx 2.1$; sources with lower values are thus likely to be RBs. Therefore, the flux ratio can serve as a diagnostic tool for distinguishing between source types when both X-ray and $\gamma$-ray fluxes are available. This represents an update to the flux ratios reported by D. Salvetti et al. (2017), in which such distinctions could not be clearly identified.
3.5.2 Optical and Infrared
[…]
Name Gaia ID sep Pan-STARRS ID sep SkyMapper ID sep 2MASS ID sep